Operation Summary for December 8, 2010 - December 10, 2010
Wednesday Dec. 8
The day shift began with Operations monitoring Tevatron store (TeV) 8340 (90.29E30) and the Recycler stash (185.5E10), with the Antiproton Source (Pbar) stacking (33.4E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.
During the day shift, Operations terminated store 8340 at 12:08 PM. Mechanical Support personnel changed some LCW filters while the TeV was at low energy; Electrical (EE) Support personnel replaced two corrector magnet power supplies as well. Operations injected a proton-only store into the TeV to get new orbit smooth values after the corrector change outs. Operations put the TeV into HEP shot setup at 3:14 PM. Unfortunately, the TeV quenched at 4 PM due to a TeV RF station (TRF4) tripping off during the ramp. Operations reported that a Recycler power supply developed a LCW leak; EE Support personnel repaired the leak.
During the evening shift, Operations began the TeV quench recovery. TeV RF personnel changed the power amplifier for RF stations TRF4 and 6. Operations prepared the CDF and D0 collision halls for accesses. Operations began the TeV checkout at 7:30 PM. Operations put the TeV into shot setup at 9:30 PM and established store 8348 at 11:35 PM with an average initial luminosity (AIL) of 198.07E30. Operations reported sending 30 minutes of beam to the returning Meson FTBF experiment T-994 (JASMIN).
Thursday Dec. 9
The midnight shift began with Operations monitoring store 8348 (144E30) and the stash (60E10), with Pbar stacking (18E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations established beam to T-944 at 4:18 AM. Operations terminated store 8348 at 4:49 AM, put the TeV into shot setup, and established store 8349 at 6:07 AM with an AIL of 184.37E30. Operations disabled beam from the Proton Source at 6:08 AM in preparation for MI and Booster accesses. The MI access started at 7 AM. The Booster access started at 7:15 AM. The NuMI access started at 7:33 AM.
During the day shift, Operations reported that the primary reason for the accesses was to find and repair a vacuum leak in the MI abort line. Operations reported that Linac, Booster, and NuMI personnel had finished their accesses by the end of shift. MiniBooNE made an access during the day as well. No beam was delivered to users due to the accesses.
During the evening shift, TeV personnel began end-of-store and crystal collimator studies at 7 PM. Vacuum technicians completed their repairs in MI at 5:50 PM. Operations established beam to MiniBooNE at 5:53 PM. Operations reported problems recovering two MI Lambertson magnets. Operations established beam to NuMI and Pbar 7:37 PM.
Friday Dec. 10
The midnight shift began with Operations monitoring store 8349 (33.61E30) and the stash (106.5E10), with Pbar stacking (24.5E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations terminated store 8349 at 1:43 AM and put the TeV into shot setup, but the shot setup was aborted at 2:30 AM when a MI bend power supply tripped off. Operators called in EE Support personnel to investigate and they decided to take the power supply out of the circuit and turn on. Operations restored beam to NuMI and Pbar at 5:54 AM. Operations put the TeV into shot setup at 6:36 AM and established store 8351 at 7:50 AM with an AIL of 170.65E30.
The Plans for Friday and the WeekendThe plans for today and the weekend are to stack, stash, and store.
Comments and Suggestions
|last modified 12/10/2010 email Fermilab|
|Security, Privacy, Legal|